# Nonisotropic Gaussian Diffusion for Realistic 3D Human Motion Prediction

### v1

@@ Abstract
Probabilistic human motion prediction aims to forecast multiple possible future movements from past observations. While current approaches report high diversity and realism, they often generate motions with undetected limb stretching and jitter. To address this, we introduce SkeletonDiffusion, a latent diffusion model that embeds an explicit inductive bias on the human body within its architecture and training. Our model is trained with a novel nonisotropic Gaussian diffusion formulation that aligns with the natural kinematic structure of the human skeleton. Results show that our approach outperforms conventional isotropic alternatives, consistently generating realistic predictions while avoiding artifacts such as limb distortion. Additionally, we identify a limitation in commonly used diversity metrics, which may inadvertently favor models that produce inconsistent limb lengths within the same sequence. SkeletonDiffusion sets a new benchmark on three real-world datasets, outperforming various baselines across multiple evaluation metrics. Visit our project page.
@@
Figure 1: SkeletonDiffusion is a latent diffusion model that takes a past motion sequence as input, and generates multiple realistic futures. We propose a novel nonisotropic diffusion formulation that focuses on the skeleton structure. Our predictions contain close to ground truth samples and also diverse but plausible alternatives. Here, we show the most diverse ensemble of three motions including the prediction closest to the ground truth among 50 generated futures.
@@ 1 Introduction
First, consider the skeleton structure and joint categories throughout the entire network and build our architecture on top of Typed-Graph Convolutions [65]. In contrast, existing SHMP approaches either ignore the skeleton’s graph structure [5, 90, 88, 16] or only leverage it at intermediate stages through Graph Convolutional Networks (GCNs)  [54, 19, 67, 80]. Second, we model the generative strategy to integrate the explicit bias. Similarly to the recent advances in SHMP based on diffusion models [5, 67, 16, 80], we opt for latent diffusion [62]. However, we replace the conventional isotropic Gaussian diffusion training [31] with a novel nonisotropic formulation that accounts for joint relations directly in the generation process: HMP is a structured problem defined by the skeleton kinematic graph, and we exploit this knowledge to define a fixed non-diagonal noise covariance for the diffusion process. To the best of our knowledge, this is the first nonisotropic diffusion process to support a non i.i.d. (i.i.d. = independent and identically distributed) latent space and reflect the problem structure. Despite demonstrating its usefulness in the skeletal domain, its applicability can be broader and touch all the domains where the conventional i.i.d. noise assumption may not hold due to the structural nature of the task.
@@
  • We propose SkeletonDiffusion, a novel latent diffusion model for SHMP that explicitly incorporates end-to-end the skeleton structure in the graph architecture and the diffusion training.
@@ 2.1 Human Motion Prediction
Probabilistic HMP has been addressed via generative adversarial networks [7, 40, 47], variational autoencoders (VAE) [77, 85, 13, 54, 54, 25, 19, 88, 82], and more recently diffusion models [67, 16, 5, 64, 80]. Among these works, HumanMAC [16] and CoMusion [67] perform diffusion in input space, relying on a transformer backbone and representing the time dimension in Discrete Cosine Space (DCT), a temporal representation widely employed in SHMP [90, 54, 19, 80]. BeLFusion [5] performs latent diffusion [62] in a semantically meaningful latent space but by leveraging a U-Net [21]. We also wish to perform diffusion in latent space, due to its speed and generalization power [10]. Differently from deterministic HMP approaches [17, 44, 43], stochastic approaches leverage Graph Convolutional Networks (GCN) [39, 74] on the skeleton graph only at intermediate stages [54, 19, 67, 80]. In our method, we build on top of Typed-Graph Convolutions [65] and design a conditional autoencoder and denoising network. We later show how this is instrumental for our novel non-isotropic formulation.
@@ 2.2 Nonisotropic Probabilistic Diffusion Models
Diffusion models [31, 62] usually specify the noisification process through isotropic Gaussian random variables, sampling the noise for each diffusion step following the i.i.d. assumption. Recent studies [24, 15] show that the isotropic noise prior may not be the best choice for all tasks: optimizing the noise at inference time may improve result quality [22] or solve related tasks [37]. Few works in image generation explore non-Gaussian or learned alternatives, motivated by their efficiency [91, 50], or by addressing image corruption and inverse problems [20, 66]. When considering nonisotropic Gaussian processes [38, 32, 33, 75], the generated images are qualitatively comparable to the isotropic ones but retain longer training and inference time and less scalability [38, 32, 33]. We present a novel nonisotropic training formulation by modifying the covariance matrix of the noise addition, making the noisification process aware of joint connections. Since we rely on the known skeleton graph, the covariance matrix is not learned [91, 50] but fixed regardless of the input motion. While covariance matrices that depend on the input might not scale well with the problem size [33], our formulation is efficient and comes at no additional computational expenses during both training and inference. To the best of our knowledge, we are the first to apply nonisotropic Gaussian diffusion to a structured problem, also showing that our formulation converges with fewer iterations and parameters than its isotropic alternative (see Sup.Mat.).
Figure 2: Our nonisotropic diffusion formulation. By diffusing a random variable ${\boldsymbol{x}}_{0}\in\mathbb{R}^{J}$ where $J$ is the number of body joints, instead of considering the joint dimensions i.i.d. as in isotropic diffusion, we take into account skeleton connections in $\boldsymbol{\Sigma}_{N}$. With the scheduler ${\gamma}_{t}$, we design a noise that transitions from isotropic to nonisotropic.
Figure 3: Overview of SkeletonDiffusion. We first train a conditional autoencoder, consisting of $\mathrm{{e}}$ and $\mathrm{{d}}$, to encode future human motions Y in a latent space $\boldsymbol{z}\in\mathbb{R}^{J\times L}$ where each of the $J$ latent joint dimensions corresponds to a human body joints. In this learned latent space, we train a denoiser network $\mathrm{{g}}$ with our novel nonisotropic diffusion formulation to generate latent embeddings ${\boldsymbol{x}}_{\theta}$ conditioned on the past X. During inference, the diffused samples are decoded to generate multiple diverse motions $\tilde{{\textbf{Y}}}$.
@@ 3 Preliminaries
3.1 Problem Formulation
The task of 3D Human Motion Prediction (HMP) takes as input a past sequence of $P$ poses and aims to predict the corresponding future as a sequence of $F$ poses. Formally, the input motion sequence is defined as $\textbf{X}=\left[{\textbf{p}}_{-P+1},{\textbf{p}}_{-P+2},\dots,{\textbf{p}}_{0}\right]\in\mathbb{R}^{P\times J\times 3}$, and the output motion sequence is defined as ${\textbf{Y}}=\left[{\textbf{p}}_{1},~{}\dots,~{}{\textbf{p}}_{F}\right]\in\mathbb{R}^{F\times J\times 3}$ with $J$ being the number of human body joints and ${\textbf{p}}_{\tau}$ the 3D body pose at timestep $\tau\in\{-P+1,-P+2,\dots,0,\dots,F\}$. Probabilistic or stochastic HMP considers a set of $N$ possible future sequences as $\tilde{{\textbf{Y}}}\in\mathbb{R}^{N\times F\times J\times 3}$ for each observation rather than a single deterministic prediction.
3.2 Isotropic Gaussian Diffusion
@@
Denoising Diffusion Probabilistic Models (DM) [31] define the forward transitions between latent random variables $p({\boldsymbol{x}}_{t}|{\boldsymbol{x}}_{t-1})$ as a linear Gaussian model $\mathcal{N}({\boldsymbol{x}}_{t};\sqrt{{\alpha}_{t}}{\boldsymbol{x}}_{t-1},{\boldsymbol{\Sigma}}_{t})$ with noise scheduler ${\alpha}_{t}$. This forward process iteratively transforms the initial variable ${\boldsymbol{x}}_{0}$ originating from the real data distribution $p({\boldsymbol{x}}_{0})$ into isotropic Gaussian noise $p({\boldsymbol{x}}_{T})=\mathcal{N}({\boldsymbol{x}}_{T};\boldsymbol{0},\mathbb{I})$. The reverse diffusion consists in sampling ${\boldsymbol{x}}_{T}\sim\mathcal{N}({\boldsymbol{x}}_{T};\boldsymbol{0},\mathbb{I})$ and iteratively applying the denoising transitions $q_{\theta}({\boldsymbol{x}}_{t-1}|{\boldsymbol{x}}_{t})$ parametrized by a neural network $\theta$ to obtain samples from the real data distribution, eventually with conditioning on other variables [5, 67] or modalities [62, 71]. These parameters are learned by maximizing the usual variational bound on log-likelihood [31, 48], derived from the closed form $p({\boldsymbol{x}}_{t}|{\boldsymbol{x}}_{0})$ of the forward transitions. Conventional isotropic diffusion defines the random variables ${\boldsymbol{x}}_{t}$ as i.i.d, i.e., as having a diagonal covariance matrix
@@
While Ho et al. [31] define the random variables in the input space, Latent Diffusion Models (LDM) [62, 71] reduce the dimensionality of the problem by considering lower-dimensional embeddings of the variables.
@@ 4 Method
We want to combine the generative capabilities of diffusion models with an explicit prior on realistic motions. This prior is obtained by embedding the knowledge about skeletal connections in both the network architecture and the training procedure. We present SkeletonDiffusion , a novel nonisotropic latent diffusion model where the noise addition is not i.i.d. but reflects limb connections. Here, we first present our nonisotropic diffusion formulation reflecting skeletal connections (Sec. 4.1) in a model-agnostic setting, and then introduce SkeletonDiffusion , implementing our nonisotropic training protocol with joint-attentive graph convolutional architecture (Sec. 4.2).
@@ 4.1 Nonisotropic Gaussian Diffusion
Clearly, in HMP every joint position depends on those of its neighbors. Relying on the i.i.d. noise assumption of conventional diffusion models [31, 62] would overlook such relations. Contrary to isotropic Gaussian diffusion that denoises all dimensions of random variables equally, we propose a nonisotropic formulation where the random variable ${\boldsymbol{x}}_{0}\in\mathbb{R}^{J}$ is denoised depending on the kinematic relations of a body with $J$ joints (Fig. 2).
@@
Since joint relationships are not dependent on the diffusion transition timestep $t$, we define our transition covariance matrix ${\boldsymbol{\Sigma}}_{t}$ in dependence of a correlation matrix $\boldsymbol{\Sigma}_{N}\in\mathbb{R}^{J\times J}$ encoding the skeleton structure:
@@
In Sup.Mat., we ablate the choice of $\mathbf{A}$ against two different, more densely populated choices (the weighted transitive closure of the skeleton graph and its masked version). Our formulation comes with negligible computational expenses without loss of generalization and can be adapted to any structural problem with a suitable choice of $\mathbf{A}$.
@@
Although the simple solution presented in Eq. 3 is already superior to the isotropic diffusion (see Sec. 5.2), we observe that different diffusion timesteps $t$ relate to different aspects of the generation process. First, the network figures out high-level, global properties of the future motion, and later, fine-grained joints’ play a more significant role. With this motivation, we define a noise addition ${\boldsymbol{\Sigma}}_{t}$ that transitions from isotropic to nonisotropic noise:
@@
where ${\gamma}_{t}=1-t/T$ defines a monotonically decreasing scheduler. See details in the Sup.Mat.
@@
To avoid computing the noise at timestep $t$ in a recursive manner and effectively be able to perform training, we derive the closed-form $p({\boldsymbol{x}}_{t}|{\boldsymbol{x}}_{0})$ of the forward process
@@
where the nonisotropic noise is obtained from isotropic noise ${\boldsymbol{\epsilon}}\sim\mathcal{N}(\boldsymbol{0},\textbf{I})$ through the Eigendecomposition of the covariance matrix ${\boldsymbol{\Sigma}}_{t}=\boldsymbol{U}{\boldsymbol{\Lambda}}_{t}\boldsymbol{U}^{\top}$, with $\bar{\boldsymbol{\Lambda}}_{t}=\tilde{\gamma}_{t}{\boldsymbol{\Lambda}}_{t}+(1-\bar{\alpha}_{t})\mathbb{I}$, and $\tilde{\gamma}_{t}=(1-{\alpha}_{t}){\gamma}_{t}+{\alpha}_{t}\tilde{\gamma}_{t-1}$. To perform inference, we define a tractable form for the posterior $q({\boldsymbol{x}}_{t-1}|{\boldsymbol{x}}_{t})$ as
$\begin{split}{\boldsymbol{x}}_{t-1}=&{\boldsymbol{\mu}}_{q}+\boldsymbol{U}{\boldsymbol{\Lambda}}_{q}{\boldsymbol{\epsilon}}\\ {\boldsymbol{\Lambda}}_{q}=&{\boldsymbol{\Lambda}}_{t}\bar{\boldsymbol{\Lambda}}_{t-1}\bar{\boldsymbol{\Lambda}}_{t}^{-1}\\ {\boldsymbol{\mu}}_{q}=&\sqrt{{\alpha}_{t}}\boldsymbol{U}\bar{\boldsymbol{\Lambda}}_{t}^{-1}\bar{\boldsymbol{\Lambda}}_{t-1}\boldsymbol{U}^{\top}{\boldsymbol{x}}_{t}\\ &+\sqrt{\bar{\alpha}_{t-1}}\boldsymbol{U}\bar{\boldsymbol{\Lambda}}_{t}^{-1}{\boldsymbol{\Lambda}}_{t}\boldsymbol{U}^{\top}{\boldsymbol{x}}_{0}\\ \end{split}$ (7)
@@
ADE $\downarrow$ FDE $\downarrow$ MMADE $\downarrow$ MMFDE $\downarrow$ APDE $\downarrow$ APD $\uparrow$ CMD $\downarrow$ str jit str jit
Alg Zero-Velocity 0.755 0.992 0.814 1.015 - 0.000 39.262 0.00 0.00 0.00 0.00
VAE TPK [77] 0.656 0.675 0.658 0.674 2.265 9.283 17.127 7.34 0.34 9.69 0.48
DLow [88] 0.590 0.612 0.618 0.617 4.243 13.170 15.185 8.41 0.40 11.06 0.58
GSPS [54] 0.563 0.613 0.609 0.633 4.678 12.465 18.404 6.65 0.29 8.98 0.37
DivSamp [19] 0.564 0.647 0.623 0.667 15.837 24.724 50.239 11.17 0.82 16.71 1.0
DM HumanMAC [16] 0.511 0.554 0.593 0.591 - 9.321 - - - - -
BeLFusion [5] 0.513 0.560 0.569 0.585 1.977 9.376 16.995 7.19 0.34 9.03 0.34
CoMusion [67] 0.494 0.547 0.469 0.466 2.328 10.848 9.636 4.04 0.25 5.63 0.52
DM SkeletonDiffusion (Ours) 0.480 0.545 0.561 0.580 2.067 9.456 11.417 3.15 0.20 4.45 0.26
Table 1: Quantitative results for AMASS dataset [52]. The best results are highlighted in bold, second-best are underlined. The symbol ‘-’ indicates that the results are not reported in the baseline work. We achieve state-of-the-art performance, while the VAE-based method with the highest diversity, DivSamp, displays the worst limb stretching and limb jitter.

### v2

@@ Abstract
Probabilistic human motion prediction aims to forecast multiple possible future movements from past observations. While current approaches report high diversity and realism, they often generate motions with undetected limb stretching and jitter. To address this, we introduce SkeletonDiffusion, a latent diffusion model that embeds an explicit inductive bias on the human body within its architecture and training. We present a new nonisotropic Gaussian diffusion formulation that aligns with the natural kinematic structure of the human skeleton and models relationships between body parts. Results show that our approach outperforms isotropic alternatives, consistently generating realistic predictions while avoiding artifacts such as limb distortion. Additionally, we identify a limitation in commonly used diversity metrics, which may favor models that produce inconsistent limb lengths within the same sequence. SkeletonDiffusion sets a new benchmark on three real-world datasets, outperforming various baselines across multiple evaluation metrics. We release the code on our project page.
@@
Figure 1: SkeletonDiffusion generates futures that are simultaneously diverse and realistic. With a nonisotropic diffusion formulation reflecting the skeleton structure, we predict motions that are plausible and semantically coherent with the input past while being highly diverse. Here, we show the most diverse ensemble of three motions including the prediction closest to the ground truth among 50 generated futures.
@@ 1 Introduction
First, we consider the skeleton structure and joint categories throughout the entire network, and build our architecture end-to-end on top of Graph Convolutional Networks (GCNs). In contrast, existing SHMP approaches either ignore the skeleton’s graph structure [5, 93, 91, 16] or only leverage it at intermediate stages [55, 19, 69, 82]. Second, we model the generative strategy to integrate the explicit bias. Similarly to the recent advances in SHMP based on diffusion models [5, 69, 16, 82], we opt for latent diffusion [64]. However, we replace the conventional isotropic Gaussian diffusion training [31] with a novel nonisotropic formulation that accounts for joint relations directly in the generation process: the HMP problem is defined by the skeleton kinematic graph, and we exploit this knowledge to define a fixed non-diagonal noise covariance for the diffusion process. To the best of our knowledge, this is the first nonisotropic diffusion process to support a non i.i.d. latent space and reflect the dependencies among components (joints) according to the given problem structure (skeleton kinematic). Despite demonstrating its usefulness in the skeletal domain, its applicability can be broader and touch all the domains where the conventional i.i.d. noise assumption may not hold.
@@
  • We propose SkeletonDiffusion, a latent diffusion model for SHMP that explicitly incorporates end-to-end the skeleton structure through the adjacency matrix in the graph architecture and the diffusion training.
@@ 2.1 Human Motion Prediction
Probabilistic HMP has been addressed via generative adversarial networks [7, 41, 48], variational autoencoders (VAE) [79, 87, 13, 55, 55, 25, 19, 91, 84], and more recently diffusion models [69, 16, 5, 66, 82]. Among these works, HumanMAC [16] and CoMusion [69] perform diffusion in input space, relying on a transformer backbone and representing the time dimension in Discrete Cosine Space (DCT), a temporal representation widely employed in SHMP [93, 55, 19, 82]. BeLFusion [5] performs latent diffusion [64] in a semantically meaningful latent space but by leveraging a U-Net [21]. We also wish to perform diffusion in latent space, due to its speed and generalization power [10]. Differently from deterministic HMP approaches [17, 45, 44], stochastic approaches leverage Graph Convolutional Networks (GCN) [40, 76] on the skeleton graph only at intermediate stages [55, 19, 69, 82]. We build on top of Typed-Graph Convolutions [67] and design a fully GCN autoencoder and denoising network, retaining the semantic meaning of body joints in latent space and thus embedding an explicit prior on the human skeletal structure in the model architecture.
@@ 2.2 Nonisotropic Probabilistic Diffusion Models
Diffusion models [31, 64] usually specify the noisification process through isotropic Gaussian random variables, sampling the noise for each diffusion step following the i.i.d. assumption. Also on manifolds [49, 89, 35], relationships between molecule components are modeled isotropically. According to recent studies [24, 15], the isotropic noise prior may not be the best choice for all tasks: optimizing the noise at inference may improve result quality [22] or solve related tasks [38]. In image generation, few explore non-Gaussian or learned alternatives, by addressing inverse problems [20, 68], or efficiency [94, 51]. When considering nonisotropic processes [39, 32, 33, 77], the generated images are qualitatively comparable but retain longer training and inference time and less scalability [39, 32, 33]. We present a novel nonisotropic training formulation by modifying the covariance matrix of the noise addition, making the noisification process aware of joint connections. Since we rely on the known skeleton graph, the covariance matrix is not learned [94, 51] but fixed regardless of the input motion. While covariance matrices that depend on the input might not scale well with the problem size [33], our formulation is efficient and comes at no additional computational expenses during both training and inference. To the best of our knowledge, we are the first to apply nonisotropic Gaussian diffusion to a structured problem, also showing that our formulation converges with fewer iterations and parameters than its isotropic alternative (see Sec. E.3).
Figure 2: Our nonisotropic diffusion formulation. By diffusing a random variable ${\boldsymbol{x}}_{0}\in\mathbb{R}^{J}$ where $J$ is the number of body joints, instead of considering the joint dimensions i.i.d. as in isotropic diffusion, we take into account skeleton connections in $\boldsymbol{\Sigma}_{N}$. With the scheduler ${\gamma}_{t}$, we design a noise that transitions from isotropic to nonisotropic. Further dimensions can be diffused isotropically.
Figure 3: Overview of SkeletonDiffusion. We first learn a latent space $\boldsymbol{z}\in\mathbb{R}^{J\times L}$ where each of the $J$ latent joint dimensions corresponds to a human body joint, by training encoder $\mathrm{{e}}$ and decoder $\mathrm{{d}}$ to reconstruct human motions sequences. Afterward, the latent joint dimension exhibits correlations similar to human body joints (Sec. 4.2). Here the denoiser network $\mathrm{{g}}$ conditioned on the past motion $\textbf{X}\in\mathbb{R}^{P\times J\times 3}$ is trained via nonisotropic diffusion (Fig. 2 and Sec. 4.1) to generate new latent codes ${\boldsymbol{x}}_{\theta}$. The generated codes are decoded into multiple diverse future motions $\tilde{\textbf{Y}}$ matching the past X and the motion closest to the GT is taken to backpropagate the training gradient.
@@ 3 Preliminaries
Problem Formulation
Human Motion Prediction (HMP) takes as input a past sequence of $P$ poses and predicts the corresponding future $F$ poses. The input motion is defined as $\textbf{X}=\left[{\textbf{p}}_{-P+1},{\textbf{p}}_{-P+2},\dots,{\textbf{p}}_{0}\right]\in\mathbb{R}^{P\times J\times 3}$, and the output as $\textbf{Y}=\left[{\textbf{p}}_{1},~{}\dots,~{}{\textbf{p}}_{F}\right]\in\mathbb{R}^{F\times J\times 3}$ with $J$ being the number of human body joints and ${\textbf{p}}_{\tau}$ the 3D body pose at timestep $\tau\in\{-P+1,-P+2,\dots,0,\dots,F\}$. Probabilistic HMP considers a set of $N$ possible future sequences as $\tilde{\textbf{Y}}\in\mathbb{R}^{N\times F\times J\times 3}$ for each observation rather than a single deterministic prediction.
Isotropic Gaussian Diffusion
@@
Denoising diffusion probabilistic models [31] define the forward transitions $p({\boldsymbol{x}}_{t}|{\boldsymbol{x}}_{t-1})$ as a linear Gaussian model $\mathcal{N}({\boldsymbol{x}}_{t};\sqrt{{\alpha}_{t}}{\boldsymbol{x}}_{t-1},{\boldsymbol{\Sigma}}_{t})$ with noise scheduler ${\alpha}_{t}$, and the random variables ${\boldsymbol{x}}_{t}$ as i.i.d with isotropic, diagonal covariance
@@
The forward process iteratively transforms the true variable ${\boldsymbol{x}}_{0}$ into isotropic Gaussian noise $p({\boldsymbol{x}}_{T})=\mathcal{N}({\boldsymbol{x}}_{T};\boldsymbol{0},\mathbb{I})$. The reverse diffusion samples ${\boldsymbol{x}}_{T}\sim\mathcal{N}({\boldsymbol{x}}_{T};\boldsymbol{0},\mathbb{I})$ and iteratively applies the denoising transitions $q_{\theta}({\boldsymbol{x}}_{t-1}|{\boldsymbol{x}}_{t})$ parametrized by a neural network $\theta$ to obtain samples from the real data distribution.
@@ 4 Method
We first present our nonisotropic diffusion formulation (Sec. 4.1), discuss its application in latent space (Sec. 4.2), and then introduce SkeletonDiffusion (Sec. 4.3).
@@ 4.1 Nonisotropic Gaussian Diffusion
Clearly, in HMP every joint position depends on those of its neighbors. Relying on the i.i.d. noise assumption of conventional diffusion models [31, 64] would overlook such relations. Contrary to isotropic Gaussian diffusion that denoises all dimensions of a random variable ${\boldsymbol{x}}_{0}\in\mathbb{R}^{J}$ equally, we propose a nonisotropic formulation where each $j$-th dimension is denoised depending on the kinematic relations of a body with $J$ joints (Fig. 2).
@@
Since joint relationships do not dependend on the diffusion timestep $t$, we define our transition covariance matrix ${\boldsymbol{\Sigma}}_{t}$ in dependence of a correlation matrix $\boldsymbol{\Sigma}_{N}\in\mathbb{R}^{J\times J}$ encoding the skeleton structure:
@@
We ablate $\mathbf{A}$ against two more sophisticated, densely populated choices (Sec. E.2). Our formulation comes with negligible computational expenses and can be adapted to any problem that can be defined by an adjacency matrix $\mathbf{A}$.
@@
Although the simple solution in Eq. 3 is already superior to isotropic diffusion (see Sec. 5.2), we observe that different diffusion timesteps $t$ relate to different aspects of the generation process. First, the network figures out high-level, global properties of the future motion, and later, fine-grained joints’ play a more significant role. With this motivation, we define a noise addition ${\boldsymbol{\Sigma}}_{t}$ that transitions from isotropic to nonisotropic noise:
@@
where ${\gamma}_{t}=1-t/T$ defines a monotonically decreasing scheduler. Detailed derivation and alternative scheduler formulation explored in early stages are in Appendix A.
@@
We derive the closed-form $p({\boldsymbol{x}}_{t}|{\boldsymbol{x}}_{0})$ of the forward process as
@@
where the nonisotropic noise is obtained from isotropic noise ${\boldsymbol{\epsilon}}\sim\mathcal{N}(\boldsymbol{0},\textbf{I})$ through the Eigendecomposition of the covariance matrix ${\boldsymbol{\Sigma}}_{t}=\boldsymbol{U}{\boldsymbol{\Lambda}}_{t}\boldsymbol{U}^{\top}$, with $\bar{\boldsymbol{\Lambda}}_{t}=\tilde{\gamma}_{t}{\boldsymbol{\Lambda}}_{t}+(1-\bar{\alpha}_{t})\mathbb{I}$, and $\tilde{\gamma}_{t}=(1-{\alpha}_{t}){\gamma}_{t}+{\alpha}_{t}\tilde{\gamma}_{t-1}$. To perform inference, we derive the tractable form for the posterior $q({\boldsymbol{x}}_{t-1}|{\boldsymbol{x}}_{t})$ as
$\begin{split}{\boldsymbol{x}}_{t-1}=&{\boldsymbol{\mu}}_{q}+\boldsymbol{U}{\boldsymbol{\Lambda}}_{q}{\boldsymbol{\epsilon}},\\ {\boldsymbol{\Lambda}}_{q}=&{\boldsymbol{\Lambda}}_{t}\bar{\boldsymbol{\Lambda}}_{t-1}\bar{\boldsymbol{\Lambda}}_{t}^{-1},\\ {\boldsymbol{\mu}}_{q}=&\sqrt{{\alpha}_{t}}\boldsymbol{U}\bar{\boldsymbol{\Lambda}}_{t}^{-1}\bar{\boldsymbol{\Lambda}}_{t-1}\boldsymbol{U}^{\top}{\boldsymbol{x}}_{t}\\ &+\sqrt{\bar{\alpha}_{t-1}}\boldsymbol{U}\bar{\boldsymbol{\Lambda}}_{t}^{-1}{\boldsymbol{\Lambda}}_{t}\boldsymbol{U}^{\top}{\boldsymbol{x}}_{0}.\\ \end{split}$ (7)
Training Objective
The KL-divergence typically employed to train denoising diffusion models [31, 73] can be formulated as Mahalanobis distance. Exploiting the Eigendecomposition, we apply the spectral theorem obtaining
$L_{\mathrm{diff}}({\boldsymbol{x}}_{\boldsymbol{\theta}},{\boldsymbol{x}}_{0},t):=\bar{\alpha}_{t}\|\bar{\boldsymbol{\Lambda}}_{t}^{-1/2}\boldsymbol{U}^{\top}({\boldsymbol{x}}_{\boldsymbol{\theta}}-{\boldsymbol{x}}_{0})\|^{2}.$ (8)
Detailed derivations are reported in Sec. A.3, together with the objective for regressing the noise [31] instead of the true variable [64]. Noticeably, as the eigenvalues ${\boldsymbol{\Lambda}}_{N}$ are fixed by the skeleton graph, all required matrices do not depend on the specific input and can be precomputed.
@@
ADE $\downarrow$ FDE $\downarrow$ MAE $\downarrow$ MMADE $\downarrow$ MMFDE $\downarrow$ APDE $\downarrow$ APD $\uparrow$ CMD $\downarrow$ str jit str jit
Alg Zero-Velocity 0.755 0.992 7.779 0.814 1.015 9.292 0.000 39.262 0.00 0.00 0.00 0.00
VAE TPK [79] 0.656 0.675 10.191 0.658 0.674 2.265 9.283 17.127 7.34 0.34 9.69 0.48
DLow [91] 0.590 0.612 8.510 0.618 0.617 4.243 13.170 15.185 8.41 0.40 11.06 0.58
GSPS [55] 0.563 0.613 9.045 0.609 0.633 4.678 12.465 18.404 6.65 0.29 8.98 0.37
DivSamp [19] 0.564 0.647 8.027 0.623 0.667 15.837 24.724 50.239 11.17 0.82 16.71 1.0
DM HumanMAC [16] 0.511 0.554 - 0.593 0.591 - 9.321 - - - - -
BeLFusion [5] 0.513 0.560 7.125 0.569 0.585 1.977 9.376 16.995 7.19 0.34 9.03 0.34
CoMusion [69] 0.494 0.547 6.715 0.469 0.466 2.328 10.848 9.636 4.04 0.25 5.63 0.52
DM SkeletonDiffusion (Ours) 0.480 0.545 6.124 0.561 0.580 2.067 9.456 11.417 3.15 0.20 4.45 0.26
Table 1: Quantitative results on AMASS [53]. The best results are highlighted in bold, second-best are underlined. The symbol ‘-’ indicates that the results are not reported in the baseline work. We achieve state-of-the-art performance, while the VAE-based method with the highest diversity, DivSamp, displays the worst limb stretching and limb jitter.
4.2 Correlated Latent Space
Extending with i.i.d Features
While our nonisotropic diffusion has been defined in Sec. 4.1 to operate on $\boldsymbol{z}\in\mathbb{R}^{J}$, here we extend the formulation to multiple dimensions and opt for a two-dimensional latent representation $\boldsymbol{z}\in\mathbb{R}^{J\times L}$, shown effective in other domains [64] but not applied to HMP before. Every $j$-th body joint is described by a feature vector of dimension $L$, dimension which does not explicitly encode information between joints. Thus we assume i.i.d noise over this dimension and diffuse it isotropically, allowing for a richer feature representation.
Correlations in Latent Space
The foundation behind our nonisotropic diffusion formulation of Sec. 4.1 is the existing correlation between body joints, described in the adjacency matrix $\mathbf{A}$ and reflected in the noisification process by the correlation matrix $\boldsymbol{\Sigma}_{N}$. In a space where no correlations exist, nonisotropic diffusion is meaningless. For this reason, in our latent space the semantic notion of body joint is intact and the correlations between each $j$-th dimension resembles human body joints motions (see Sec. E.1 and Fig. 9).